Public health and safety officials have classified the Zika virus as a hazardous threat to pregnant women in Brazil. A study based in Rio found the virus to be more active near a body of water that was found to be polluted. A study of 20 water samples found an average chemical waste of 168 g per mile with a standard deviation of 17.81 g.


Find the 95% confidence interval (in g per mile) for this study. (Use a table or technology. Round your answers to three decimal places.)

(. , ) g per mile
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 168

sample standard deviation = s = 17.81

sample size = n = 20

Degrees of freedom = df = n - 1 = 20 - 1 = 19

At 95% confidence level

= 1 - 95%

=1 - 0.95 =0.05

/2 = 0.025

t/2,df = t0.025,19 = 2.093

Margin of error = E = t/2,df * (s /n)

= 2.093 * (17.81 / 20)

Margin of error = E = 8.335

The 95% confidence interval estimate of the population mean is,

  ± E

= 168 ± 8.335

= ( 159.665, 176.335 )